### v1

# GraphAU-Pain: Graph-based Action Unit Representation for Pain Intensity Estimation

Abstract
Understanding pain-related facial behaviors is essential for digital healthcare in terms of effective monitoring, assisted diagnostics, and treatment planning, particularly for patients unable to communicate verbally. Existing data-driven methods of detecting pain from facial expressions are limited due to interpretability and severity quantification. To this end, we propose GraphAU-Pain, leveraging a graph-based framework to model facial Action Units (AUs) and their interrelationships for pain intensity estimation. AUs are represented as graph nodes, with co-occurrence relationships as edges, enabling a more expressive depiction of pain-related facial behaviors. By utilizing a relational graph neural network, our framework offers improved interpretability and significant performance gains. Experiments conducted on the publicly available UNBC dataset demonstrate the effectiveness of the GraphAU-Pain, achieving an F1-score of $66.21\%$ and accuracy of $87.61\%$ in pain intensity estimation. The code is available for re-implementation at github.com/ZW471/GraphAU-Pain.
I INTRODUCTION
Pain detection is critical in clinical and caregiving settings for timely assessment and improved patient management. Current methods like self-reports and observational evaluations have notable limitations [10]. Self-reports are subjective and depend on patient communication ability, often impaired in nonverbal individuals, children, or those with cognitive impairments. Observational methods, while more objective, require extensive training to ensure accuracy and consistency.
Deep learning has driven interest in automated pain estimation via facial expression analysis, bypassing advanced medical equipment while providing objective measures. Methods like CNNs [22, 24] and hybrid frameworks [7, 29, 1] have been applied for pain estimation using facial features. Transformers [13, 28] perform remarkably in pain prediction from facial videos. However, most rely solely on image features, neglecting physiological insights, limiting clinical interpretability. Additionally, undersampling to address dataset imbalance reduces generalizability in diverse populations [2, 3].
Pain estimation can leverage features tied to facial expressions. Landmark-based methods like nose tip or eye corner coordinates outperform pixel-based techniques but lack clear physiological links to pain [8, 12]. The Facial Action Coding System (FACS) [6] maps facial movements into Action Units (AUs) with intensity levels, aggregating into the Prkachin and Solomon Pain Intensity (PSPI) score [23]. The UNBC dataset [19] provides AU and PSPI-labeled videos, supporting AU-informed methods like K-Nearest Neighbor [30] and Bayesian Networks [9], which report high accuracy but suffer from overoptimism due to class imbalance and overlook AU relationships. Recent approaches like GLA-CNN [26] and Multiple Instance Learning [4] have improved AU relation modeling. However, these methods rely on pixel-wise AU relationships, struggling with subtle facial changes and generalization across diverse racial groups, limiting real-world applicability.
Recent advances in graph neural networks (GNNs), such as Multi-dimensional Edge Feature-based AU Relation Graph for AU (ME-GraphAU) [20] and Graph Relation Network (GRU) [25], have demonstrated promising AU prediction performance on the DISFA dataset [21] and BP4D dataset [31]. These methods use CNN backbones like ResNet [11] or VGG [14] to learn image features for each AU, then construct a relational graph based on the AU features. Each node in such graphs represents an AU, and each edge represents the relationship between a pair of AUs [17]. Through message propagation in GNN layers, the output AU features can capture individual AU information from neighbors as well as structural information [15]. Such graph-based AU features can then be aggregated to construct a full-face representation informed by AUs and their relationships for downstream tasks [16], inspiring us to incorporate graph-based AU detection into pain intensity estimation.
Motivated by the above, we introduce GraphAU-Pain and summarize our three contributions as follows:
  • GraphAU-Pain Model. We fuse a ResNet-50 backbone with graph-based AU representation learning, addressing the shortcomings of purely image-driven methods.
  • Cross-Dataset Transfer with Relabeling. To mitigate poor direct training on UNBC, we fine-tune GraphAU-Pain on DISFA-pretrained weights (sharing most AUs), partially relabeling UNBC to create UNBC+ to handle missing AU annotations and boost performance.
  • High Performance & Interpretability. Comprehensive experiments validate GraphAU-Pain for pain estimation, outperforming GLA-CNN [26] (current state-of-the-art work in this area) and demonstrating improved interpretability via explicit AU modeling.
II MATERIALS AND METHODS
II-A Implementation Overview
The GraphAU-Pain model is designed to estimate pain intensity based on facial image data, and its training involves two key steps. First, the full-face and AU representation learning modules are trained for AU occurrence prediction. These modules are adapted from the AU Relationship-aware Node Feature Learning (ANFL) component of ME-GraphAU [20]. With only a $20\%$ average F1-score achieved when directly training ANFL for AU occurrence prediction on the UNBC dataset, we developed a cross-dataset transfer learning strategy to improve performance. This strategy transfers AU prediction capabilities from the DISFA dataset model trained by Luo et al. [20] to the UNBC dataset. To address differences in AU labels between datasets, we created a relabeled UNBC+ dataset to align the AU annotations and used undersampling to address AU imbalance. Second, with the weights trained for AU prediction, GraphAU-Pain is then trained for pain intensity estimation with the full UNBC dataset. Facial images are first processed through the CNN backbone to extract pixel-wise full-face representations. These representations are then transformed into graph-based structures to learn AU-specific features. Finally, the full-face and AU-based features are combined to estimate pain intensity. The performance of the model is evaluated using two metrics: average F1-score and accuracy.
As illustrated in Fig. 1, the GraphAU-Pain model comprises three sequential modules: Full-face Representation Learning that extracts high-level facial features through pixel-based global representations, AU Representation Learning that captures both local and global AU using graph-based representations, and a Pain Intensity Classifier that maps the learned features to specific pain intensity levels.
Figure 1: Illustration of the GraphAU-Pain model architecture.
A global CNN backbone (orange) extracts a full-face representation, which is split into three feature branches by the AU representation learning module (green) and processed by the pain level classification module (blue). ANFL denotes the AU Relationship-aware Node Feature Learning component, FC the fully-connected layer, and MatMult matrix multiplication.
II-B Fine-tuning Representation Learning Modules
Our preliminary experimental results showed that directly training ANFL on the UNBC dataset failed to provide accurate AU occurrence prediction. Similarly, directly training GraphAU-Pain on UNBC yielded unsatisfactory results, with $30\sim 40\%$ F1-score and $60\sim 80\%$ accuracy after trying several settings. To tackle this problem and improve AU prediction performance through transfer learning, in this work we fine-tuned ANFL’s weights pretrained on DISFA, which shares 6 out of 8 AUs with UNBC, 3 of them used in PSPI calculation. To relabel UNBC so its AU labels align with DISFA, the facial images of UNBC are passed to the pretrained ANFL module to obtain 8 predicted AU labels. Of these 8 labels, the 6 that exist in UNBC are replaced with the original values, while for AU1 and AU2 predicted values are used. We refer to this relabelled dataset as UNBC+ to distinguish it from the original UNBC dataset.
II-C Model
Full-face Representation Learning
GraphAU-Pain uses a ResNet-50 backbone for extracting a full-face representation from an input image. By inputting a face image $\mathbf{x}\in\mathbb{R}^{172\times 172\times 3}$ to the backbone, we obtain $\mathbf{h_{b}}\in\mathbb{R}^{36\times 2048}$, which represents 36 facial image features of length 2048, each corresponding to a position in the image.
AU Representation Learning
After acquiring the backbone feature $\mathbf{h_{b}}$, the AU representation module is applied to learn two distinct embeddings: $\mathbf{h_{a}}\in\mathbb{R}^{d_{\text{AU}}}$, where $d_{\text{AU}}=8$, encoding each AU’s occurrence; and $\mathbf{h_{g}}\in\mathbb{R}^{d_{\text{node}}}$, where $d_{\text{node}}=512$, encoding the complete structure of the AU relational graph. The AU representation module first learns the AU relational graph $\mathcal{G}$ from $\mathbf{h_{b}}$ as in the original ANFL. It begins with applying $d_{\text{AU}}$ fully connected layers on $\mathbf{h_{b}}$ to obtain a set of initial AU representations $\mathbf{V}\in\mathbb{R}^{d_{\text{AU}}\times d_{\text{node}}}$ as node features. Each node then connects with its $K=3$ most similar nodes based on dot-product similarity. Next, messages are passed through a graph convolutional layer:
$\mathbf{V}^{\prime}\;=\;\text{ReLU}\Bigl{[}\mathbf{V}+\text{BN}\bigl{(}\mathbf{A}^{\top}\text{FC}_{1}(\mathbf{V})+\text{FC}_{2}(\mathbf{V})\bigr{)}\Bigr{]}\;\in\;\mathbb{R}^{d_{\text{AU}}\times d_{\text{node}}},$ (1)
where BN refers to batch normalization, FC refers to fully connected layers, and $A$ is the normalized adjacency matrix. Then, AU features $\mathbf{h_{a}}\in\mathbb{R}^{d_{\text{AU}}}$ are obtained by a similarity-calculating strategy [20], and the graph representation $\mathbf{h_{g}}\in\mathbb{R}^{d_{\text{node}}}$ is computed by a global sum pooling of the node embeddings. Although it would be possible to add an edge feature learning module here [20], we omitted it to avoid overfitting on the UNBC dataset. After obtaining these three features, three fully connected layers with ReLU activation map each of ${\mathbf{h_{b}}}$, ${\mathbf{h_{a}}}$, and ${\mathbf{h_{g}}}$ to a common dimension of $36$, producing ${\mathbf{h_{b}^{\prime}}}$, ${\mathbf{h_{a}^{\prime}}}$, and ${\mathbf{h_{g}^{\prime}}}$. (For $\mathbf{h_{b}}$, FC is applied row-wise.) Finally, a feature infusing step on $\mathbf{h_{a}}^{\prime}$ and $\mathbf{h_{b}}^{\prime}$ is performed:
$\mathbf{h_{ab}}\;=\;\text{ReLU}(\mathbf{h_{a}^{\prime\top}}\,\mathbf{h_{b}}^{\prime})\;\in\;\mathbb{R}^{36}.$ (2)
Representation Classifier
The final pain intensity classification is get by concatenating the interaction scalar $\mathbf{h_{ab}}$ with the $\mathbf{h_{g}}^{\prime}$ feature, then passing the result to an FC layer:
$\mathbf{\hat{y}}\;=\;\mathbf{W}_{y}\,[\mathbf{h_{ab}}\parallel\mathbf{h_{g}^{\prime}}]+\mathbf{b}_{y}\;\in\;\mathbb{R}^{d_{\text{pain}}},$ (3)
where $d_{\text{pain}}=3$ for the one-hot encoding of the three-level pain intensity classification used in this work.
II-D Loss Function
With over 80% of the frames demonstrating no expression of pain, the class imbalance in the UNBC dataset poses a significant challenge for deep learning models. To minimize this, we employ a weighted cross-entropy loss to prioritize underrepresented classes. It is calculated by
$\mathcal{L}=-\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{C}w_{j}\,y_{i,j}\,\log\max\bigl{(}p_{i,j},\epsilon\bigr{)},$ (4)
where $N$ is the number of samples, $C$ is the number of classes, $p_{i,j}$ are the softmax output probabilities for the $i$-th sample belonging to the $j$-th class, $y_{i,j}$ is a binary indicator (1 if the $i$-th sample belongs to the $j$-th class, otherwise 0), $w_{j}$ are class weights, and $\epsilon=1e^{-8}$ prevents $\log(0)$. The weight of a class $c_{j}$ is calculated by
$w_{j}=C\cdot\frac{1/\text{occurrence\_rate}(c_{j})}{\sum_{k=1}^{C}1/\text{occurrence\_rate}(c_{k})}.$ (5)
In this work, the class weights are $0.07$ for No Pain, $0.33$ for Mild Pain, and $2.6$ for Obvious Pain.
III EXPERIMENTS
III-A Datasets and Labels
GraphAU-Pain was trained on the UNBC-McMaster Shoulder Pain Expression Archive Database [19]. The dataset contains 48,398 colored frames from 25 participants with shoulder problems, showing facial expressions during pain-inducing actions. The faces were detected with OpenCV’s haarcascade frontalface default classifier and cropped to $172\times 172$. Each frame has 10 AU intensities ($0-5$) and PSPI scores ($0-16$), calculated as $PSPI=AU4+\max(AU6,AU7)+\max(AU9,AU10)+AU43$ [23]. PSPI pain intensity is categorized into ordinal levels: No Pain ($PSPI=0$), Mild Pain ($PSPI\in[1,4]$), and Obvious Pain ($PSPI\geq 5$). The categories are distributed in a skewed way, with each category respectively consisting approximately $82\%$, $15\%$ and $3\%$ of the UNBC dataset. This categorization is more clinically meaningful and interpretable compared to the method used in the GLA-CNN paper [26], as it mitigates the high sensitivity of PSPI scores to AU variations. Moreover, we replaced AU intensity with occurrence by capping the AU score at one.
III-B Training and Evaluation Details
To prepare AU occurrence prediction before pain intensity estimation training, the supervised fine-tuning (SFT) process of GraphAU-Pain’s AU representation learning module was performed on the ANFL component with weights pretrained on DISFA for 20 epochs provided by Luo et al. [20]. Our relabeled UNBC+ dataset was used in the SFT process and undersampled to address data imbalance. The undersampling process involved randomly removing approximately $90\%$ of facial images with $PSPI=0$ and excluding facial images without active AUs. The full list of frames included in this subset is made available in the code repository. Through SFT for 17 epochs with a learning rate of $1e^{-5}$, a batch size of 16, and an Adam optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.999$, and a weight decay of $5e^{-4}$, the AU representation learning module achieved remarkable results in AU detection, compared to state-of-the-art results, as detailed in Table I. The actual training for pain intensity estimation used the same hyperparameters as the SFT process but was performed on the full original UNBC dataset.
TABLE I: F1-scores and accuracies achieved for each AU by the fine-tuned ANFL module.
AU1 AU2 AU4 AU6 AU9 AU12 AU25 AU26 Avg
F1 63.09 88.65 70.74 88.46 31.94 92.12 63.93 6.63 63.20
Acc. 90.38 79.88 92.04 79.62 91.43 86.09 83.11 85.21 85.98
The GraphAU-Pain model was trained using a learning rate of $1e^{-4}$, a batch size of 64, and an Adam optimizer with the same hyperparameters used in SFT. The representation learning module is set to connect edges between an AU node and its 3 most similar AU nodes. The weights learned through SFT on ANFL were used to initialize the ResNet backbone and the representation learning module of GraphAU-Pain. The model was trained on the full UNBC dataset for 8 epochs on an NVIDIA GeForce RTX 4070 GPU (8 GB) and an Intel i7-13900H CPU, with an estimated training time of about 3 minutes per epoch. The evaluation uses the same metrics as the state-of-the-art method GLA-CNN [26]: accuracy, average F1, average recall, and average precision, where all average values are unweighted.
TABLE II: GraphAU-Pain’s performance statistics.
Metric No Pain Mild Obvious Overall
F1-score 93.10 51.19 54.35 66.21
Recall 92.85 52.48 51.02 65.45
Precision 93.35 49.96 58.14 67.15
Accuracy - 87.61
III-C Model Performance
Overall, GraphAU-Pain achieves a commendable average F1-score of $66.21\%$ and a high accuracy of $87.61\%$. Table II details the per-class results, showing strong performance for the No Pain category, which has an F1-score of $93.10\%$. However, performance declines for the Mild and Obvious categories, reflecting F1-scores of $51.19\%$ and $54.35\%$, respectively. This substantial difference is largely attributable to the dataset’s pronounced class imbalance. The log-scaled confusion matrix in Figure 2 further illustrates the distribution of predictions. While No Pain dominates the diagonal, indicating high accuracy there, some off-diagonal misclassifications occur between Mild and Obvious, and there is a noticeable bias toward predicting No Pain. Consequently, while the model is robust in detecting No Pain, additional strategies are needed to better distinguish between higher pain intensities.
(a) Classifying 3 Pain Intensities
(b) Classifying 4 Pain Intensities
Figure 2: Confusion matrix of GraphAU-Pain
The numbers are transformed in $\log_{10}$ scale for better visualization; the original number is shown below the scaled one.
III-D Comparison to SOTA
To the best of our knowledge, GLA-CNN [26] is the only other method that uses AUs for pain-intensity estimation on UNBC while focusing on cross-sectional facial image frames. No additional methods apply exactly the same categorization scheme, so we compare GraphAU-Pain with GLA-CNN and other models reported in [26]. To align labels, we reclassify pain intensities into four categories: No Pain ($PSPI=0$), Weak Pain ($PSPI=1$), Mild Pain ($PSPI=2$), and Strong Pain ($PSPI\geq 3$). By altering only the model’s final layer from 3 to 4 outputs and keeping other settings unchanged, GraphAU-Pain shows substantial gains in both accuracy and average F1-score, as shown in Table III and Figure 2(b). Note that GLA-CNN and the other compared models were trained and evaluated on an undersampled subset of UNBC [26] to deal with class imbalance, whereas GraphAU-Pain is trained on the full dataset. Therefore, their published performance might be higher than what would have been achieved on the full dataset.
TABLE III: Comparing classification performance of GraphAU-Pain with other models.
The last 3 models are retrained and tested on the same dataset as GLA-CNN, but only accuracy is reported [26].
Model Accuracy F1-score Recall Precision
GraphAU-Pain 82.40 43.83 52.96 39.96
GLA-CNN [26] 56.45 36.52 34.08 43.23
MA-NET [32] 55.16 - - -
Swin Transformer [18] 54.40 - - -
Attention CNN [27] 51.10 - - -
III-E Ablation Analysis
The ablation analysis in Table IV underscores the critical role of graph representation and GNN in the GraphAU-Pain model. Removing the graph representation (w/o graph rep.) reduces the mean F1-score from $66.2\%$ to $63.1\%$, mainly due to the performance drops in No Pain and Mild Pain, highlighting the value of graph modeling for capturing AU relationships. Similarly, removing the GNN layer (w/o GNN) causes a significant drop to $40.3\%$, emphasizing the importance of graph-based interactions in learning AU features. The simplest setup (Only ResNet), relying solely on CNNs, achieves the lowest mean F1-score of $35.2\%$, demonstrating that CNNs alone fail to effectively model AU correlations for pain estimation. These results affirm the superiority of graph-based learning methods for pain estimation.
TABLE IV: Ablation analysis for GraphAU-Pain.
w/o graph rep. removes the graph representation from the model. w/o GNN removes the GNN layer. Only ResNet relies solely on CNNs. The highest value for each category and metric is bold; the second highest is underlined.
Model Metric No Pain Mild Obvious Mean
Full F1 93.1 51.2 54.3 66.2
Recall 92.9 52.5 51.0 65.4
Precision 93.4 50.0 58.1 67.2
w/o graph rep. F1 87.7 46.4 55.2 63.1
Recall 80.8 73.4 51.0 68.4
Precision 95.9 33.9 60.2 63.4
w/o GNN F1 86.4 30.7 4.0 40.3
Recall 79.9 43.2 10.2 44.4
Precision 93.9 23.8 2.5 40.1
Only ResNet F1 70.3 31.0 4.2 35.2
Recall 56.5 68.3 15.3 46.7
Precision 92.8 20.1 2.5 38.5
III-F Limitations
While the proposed method advances pain estimation, it has key limitations. Firstly, since PSPI only captures facial expressions, it may not reflect true subjective pain [5]. Future research could focus on finding alternative pain indicators. Secondly, aligning UNBC with DISFA through UNBC+ removes three pain-related AUs and adds noise through predicted labels, potentially impacting performance. A promising direction is to design AU models specifically for pain-oriented datasets like UNBC. Lastly, although the AU occurrence-based ANFL module provides satisfactory representation, AU intensity-based approaches (e.g., GRN) could be explored, as AU intensity better relates to pain intensity.
IV CONCLUSION
This paper presents GraphAU-Pain, a GNN-based model combining graph-based AU features with full-face representation for pain prediction. It surpasses the state-of-the-art methods while enabling AU-informed pain estimation for clinical transparency. GraphAU-Pain addresses challenges like limited data and class imbalance in the UNBC dataset. Key contributions include improved pain classification benchmarks, better interpretability through AU-based representations, and critical baselines for future AU-based pain estimation. Overall, this work demonstrates the potential of GNNs for accurate, clinically viable pain estimation solutions.
ETHICAL IMPACT STATEMENT
This study involved secondary analyses of an existing dataset. No human subject data are collected in this study. All experiments and analysis are conducted on a publicly available dataset that has been described and cited accordingly. Implementation code are open source, including preprocessing, model training, and post-hoc analysis, via github.com/ZW471/GraphAU-Pain.